# Towards Heterogeneous Continual Graph Learning via Meta-knowledge Distillation

### v1

@@ 4.1 Fast Adaptation with Gradient-based Meta-Learning
4.2 Efficient Heterogeneous Subgraph Sampling
In addition to adapting to new data patterns, overcoming catastrophic forgetting is crucial, as previously observed patterns may reappear in practical scenarios. While many existing experience replay methods primarily consider homogeneous graphs or single-type nodes, they fail to capture the complex semantic and structural dependencies in heterogeneous graphs. To address this, we propose an efficient subgraph sampling strategy explicitly tailored for heterogeneous graphs, which jointly considers the diversity of nodes within the target node type and the structural connections with nodes of other types. Our method proceeds in two stages: (1) selecting representative target-type nodes by maximizing diversity in the feature or embedding space, and (2) expanding to other types of nodes through relation-type-aware importance estimation, ensuring that both semantic diversity and heterogeneous structural patterns are preserved.
Step 1: Target Node Selection via Coverage Maximization. Prior research has shown that different data points contribute unequally to model training: some significantly enhance performance, while others offer limited benefit toneva2018empirical . Motivated by this, we identify key nodes from the current task by maximizing node diversity using the Coverage Maximization (CM) strategy. This approach allows us to retain essential historical information with only a small number of nodes. Formally, given the training node set $V_{t}^{tr}$ of task $\mathcal{T}_{t}$, CM selects a subset by maximizing the coverage of the attribute/embedding space:
$\mathcal{N}(v_{i})=\{v_{j}\mid\text{dist}(v_{i},v_{j})<d,\,\mathcal{Y}(v_{i})\neq\mathcal{Y}(v_{j})\},$ (2)
@@ 4.2 Efficient Heterogeneous Subgraph Sampling
$\mathcal{C}_{r}=\bigcup_{v\in V_{\tau_{t}}}\left\{u\,\middle|\,(v,u)\in E_{r}\lor(u,v)\in E_{r}\right\},$ (3)
@@
$V^{*}_{\tau}=\text{topk}_{v\in\mathcal{C}_{\tau}}\left(\pi(v),|\mathcal{B}_{\tau}|\right),$ (5)
@@
$V^{\prime}=V_{\tau_{t}}\cup\bigcup_{\tau\neq\tau_{t}}V_{\tau},\quad E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\},$ (6)
@@
$\mathcal{L}$ $=\mathcal{L}_{\mathcal{T}_{t}}(\mathcal{G}_{t},\theta)+\lambda_{er}\mathcal{L}_{er}(\mathcal{G}_{sub},\theta)$ (7)
@@
where $\hat{\mathcal{Y}}(v_{i})$ is the predicted label of node $v_{i}$, $\lambda_{er}$ $\lambda_{er}$ modulates the contribution of buffered data in the overall loss.
4.3 Heterogeneity-aware Knowledge Distillation for Task Alignment
@@
Logit-level Distillation Inspired by previous distillation-based approaches tian2023knowledge , we transfer the soft knowledge from the teacher model (previous task classifier) to the student model (current task classifier), enabling the student model to learn both new knowledge from the current task and retained knowledge from past tasks. Mimicking teacher’s prediction results enables the student model to learn the secondary information from previous tasks that cannot be expressed by the experience replay data alone. Soft knowledge from teacher model is formulated as the predicted probability of the labels in the current task data:
$P^{T}(z_{i},t)=\text{Softmax}(f_{T}(z_{i}),t)=\frac{\exp\left[f_{T}(z_{i})/t\right]}{\sum_{j}\exp\left[f_{T}(z_{j})/t\right]},$ (8)
where $z_{i}$ is the embedding of node $v_{i}$ in $\mathcal{V}_{i}^{tr}$, $f_{T}(z_{i})$ is the score logit of $z_{i}$ obtained from teacher model, and $t$ is the temperature index to soften the peaky softmax distribution hinton2015distilling . Thus, the knowledge distillation loss of the teacher model and the student model is defined as follows:
$\mathcal{L}_{\text{logit}}=\text{Mean}(t^{2}\sum_{l}^{N}\sum_{v_{i}\in\mathcal{V}_{i}^{tr}}P_{l}^{T}(z_{i},t)\log\frac{P_{l}^{T}(z_{i},t)}{P^{S}(z_{i},t)}),$ (9)
@@
$\mathcal{L}_{\text{sem}}=\sum_{m=1}^{M}\|\boldsymbol{\alpha}_{P_{m}}^{(T)}-\boldsymbol{\alpha}_{P_{m}}^{(S)}\|_{2}$ (10)
where $\boldsymbol{\alpha}_{P_{m}}\in\mathbb{R}^{|V^{tr}|}$ is the normalized attention vector over all nodes for metapath ${P_{m}}$. We provide a detailed discussion in Appendix A.2 on how this module can be extended to arbitrary HGNNs beyond metapath-based architectures.
@@
$\mathcal{L}_{kd}=\lambda_{\text{logit}}\mathcal{L}_{\text{logit}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}},$ (11)
@@
$\mathcal{L}_{joint}=\mathcal{L}_{\mathcal{T}_{t}}+\lambda_{er}\mathcal{L}_{er}+\lambda_{kd}\mathcal{L}_{kd},$ (12)
@@
$\theta^{*}=\theta-\beta\nabla_{\theta}\mathcal{L}_{joint}(\theta),$ (13)
where $\beta$ is learning rate. By minimizing $\mathcal{L}_{joint}$, parameters $\theta$ of HGNN model are optimized for downstream node classification task. The detailed framework of MKD is provided in Appendix 1.
5 Experiments
We conducted experiments on four datasets: DBLP lv2021we , IMDB lv2021we , Freebase lv2021we and Yelp yang2020heterogeneous to evaluate the performance of MKD. We adopt three widely used HGNN backbones in our experiments: HAN wang2019heterogeneous , MAGNN fu2020magnn , and HGT hu2020heterogeneous . See Appendix  B.1 for the details of the datasets and  B.4 for experiment setup.
Baselines To evaluate the effectiveness of our proposed method, we select strong baselines from the Continual Graph Learning Benchmark (CGLB) zhang2022cglb , including Elastic Weight Consolidation (EWC) kirkpatrick2017overcoming , Memory Aware Synapses (MAS) Aljundi_2018_ECCV , Experience Replay GNN (ER-GNN) zhou2021overcoming , and Topology-aware Weight Preserving (TWP) liu2021overcoming . EWC and MAS were previously not designed for graphs, so we train a HGNN on new tasks but ignore the graph structure when applying continual learning strategies. Additionally, we include two experience-replay-based baselines: MetaCLGraph unal2023meta and FTF-ER pang2024ftf . Furthermore, we use Finetune method (without any continual learning techniques) as a lower bound, and Joint Train (which allows access to previous data during training) as an approximate upper bound caruana1997multitask .
Metrics In our experiments, Average Performance (AP) and Average Forgetting (AF) lopez2017gradient , are used to measure the performance on test sets. AP and AF are defined as $\text{AP}=\frac{1}{T}\sum_{t=1}^{T}a_{T,t},\,\,\,\,\text{AF}=\frac{1}{T-1}\sum_{t=1}^{T-1}\left(a_{T,t}-a_{t,t}\right),$ where $T$ is the total number of tasks and ${a_{i,j}}$ is the accuracy of the model on the test set of task $j$ after it is trained on task $i$.
Table 1: Performance comparison with different HGNN backbones on three benchmark datasets. The symbol $\uparrow$ ($\downarrow$) means higher (lower) is better. The best results are highlighted in bold, while the second best results are underlined. "OOM" means that the model runs out of memory on large graphs.
@@ 5 Experiments
EWC 85.4 $\pm$ 5.5 21.5 $\pm$ 10.6 69.6 $\pm$ 1.2 25.8 $\pm$ 2.4 58.3 $\pm$ 2.4 18.4 $\pm$ 4.2
MAS 91.1 $\pm$ 0.8 8.3 $\pm$ 1.5 72.0 $\pm$ 2.3 20.5 $\pm$ 3.9 59.0 $\pm$ 1.0 14.8 $\pm$ 0.9
TWP 90.5 $\pm$ 1.3 10.1 $\pm$ 2.6 74.8 $\pm$ 2.7 14.5 $\pm$ 4.4 60.9 $\pm$ 4.3 10.8 $\pm$ 6.9
ER-GNN 89.7 $\pm$ 2.9 13.1 $\pm$ 6.1 74.9 $\pm$ 2.3 12.7 $\pm$ 4.2 56.5 $\pm$ 1.3 19.6 $\pm$ 0.9
FTF-ER 90.8 $\pm$ 1.4 9.8 $\pm$ 2.4 72.0 $\pm$ 4.2 19.7 $\pm$ 6.4 58.1 $\pm$ 0.7 14.6 $\pm$ 0.6
MetaCLGraph 89.9 $\pm$ 0.9 11.4 $\pm$ 2.3 74.8 $\pm$ 3.2 14.8 $\pm$ 5.7 59.4 $\pm$ 2.5 13.0 $\pm$ 4.1
@@
EWC 89.1 $\pm$ 1.2 10.7 $\pm$ 1.8 77.3 $\pm$ 0.4 12.5 $\pm$ 0.2 69.6 $\pm$ 3.0 14.7 $\pm$ 2.9
MAS 89.8 $\pm$ 2.5 9.0 $\pm$ 3.0 77.5 $\pm$ 1.2 11.8 $\pm$ 2.6 70.4 $\pm$ 3.1 14.8 $\pm$ 2.0
TWP 89.1 $\pm$ 2.1 10.3 $\pm$ 1.1 77.7 $\pm$ 0.6 11.1 $\pm$ 1.6 69.6 $\pm$ 2.0 15.1 $\pm$ 0.9
ER-GNN 90.7 $\pm$ 1.5 3.5 $\pm$ 0.4 77.3 $\pm$ 2.2 9.0 $\pm$ 3.6 69.5 $\pm$ 1.4 13.6 $\pm$ 1.7
MetaCLGraph 90.0 $\pm$ 0.7 10.9 $\pm$ 1.4 76.1 $\pm$ 0.3 12.2 $\pm$ 1.8 69.1 $\pm$ 2.2 16.3 $\pm$ 2.5
FTF-ER 89.4 $\pm$ 2.2 5.5 $\pm$ 1.9 76.1 $\pm$ 1.4 8.9 $\pm$ 3.5 69.7 $\pm$ 2.8 13.0 $\pm$ 2.5
@@
EWC 91.0 $\pm$ 2.3 10.9 $\pm$ 4.5 73.4 $\pm$ 1.6 16.7 $\pm$ 2.5 OOM OOM
MAS 92.1 $\pm$ 2.1 7.8 $\pm$ 5.3 74.3 $\pm$ 2.0 14.2 $\pm$ 3.7 OOM OOM
TWP 91.6 $\pm$ 2.1 7.4 $\pm$ 5.0 74.0 $\pm$ 0.8 13.7 $\pm$ 3.3 OOM OOM
ER-GNN 90.0 $\pm$ 2.1 12.4 $\pm$ 4.3 75.3 $\pm$ 0.9 10.3 $\pm$ 1.5 OOM OOM
FTF-ER 92.8 $\pm$ 0.8 1.9 $\pm$ 2.3 73.0 $\pm$ 0.6 15.8 $\pm$ 1.4 OOM OOM
MetaCLGraph 92.3 $\pm$ 0.4 2.1 $\pm$ 0.7 75.6 $\pm$ 1.0 11.8 $\pm$ 1.3 OOM OOM
@@
Performance Evaluation We first systematically evaluate the extent of catastrophic forgetting in three HGNNs. From Table 1, we can observe that all HGNNs exhibit catastrophic forgetting on previous tasks. For example, on the DBLP dataset, the average forgetting (AF) for HAN and MAGNN is over 26% and 34%, respectively; on the IMDB dataset, the AF for all HGNNs is over 26%, 17%, and 21%, respectively.
We evaluate the performance of MKD and other baselines on three datasets and three backbone HGNN models. We report only the results of HAN on the Yelp dataset in Table 4, as the other two models run out of memory on this dataset. Additionally, we adapt both 2-way (i.e., two node classes per task) and 3-way settings for the Yelp dataset. We report the average values and the standard deviations over 5 runs. It shows the proposed MKD achieves superior performance across all datasets and HGNN backbones. It is noteworthy that, in some cases, FTF-ER, MetaCLGraph, and ER-GNN perform better than MKD in terms of AF, despite their lower AP scores compared to MKD. This is because these methods rely on retraining nodes from previously learned tasks. Once enough nodes are sampled, these models can largely mitigate catastrophic forgetting. However, this also sacrifices performance when learning new tasks. Although these methods attempt to address the new-old task trade-off through different strategies (e.g., ER-GNN reweights the loss between new and old tasks), they still cannot fully avoid the performance drop. We further discuss this in Appendix C.3.

### v2

@@ 4.1. Fast Adaptation with Gradient-based Meta-Learning
4.2. Diversity Sampling with semantic Consistency
In addition to adapting to new data patterns, overcoming catastrophic forgetting is crucial, as previously observed patterns may reappear in practical scenarios. While many existing experience replay methods primarily consider homogeneous graphs or single-type nodes, they fail to capture the complex semantic and structural dependencies in heterogeneous graphs. To address this, we propose DiSCo (Diversity Sampling with semantic Consistency), which jointly considers the diversity of nodes within the target node type and the structural connections with nodes of other types. DisCo proceeds in two stages: (1) selecting representative target-type nodes by maximizing diversity in the feature or embedding space, and (2) expanding to other types of nodes through relation-type-aware importance estimation, ensuring that both semantic diversity and heterogeneous structural patterns are preserved.
Step 1: Target Node Selection via Coverage Maximization. Prior research has shown that different data points contribute unequally to model training: some significantly enhance performance, while others offer limited benefit (Toneva et al., 2018). Motivated by this, we identify key nodes from the current task by maximizing node diversity using the Coverage Maximization (CM) strategy. This approach allows us to retain essential historical information with only a small number of nodes. Formally, given the training node set $V_{t}^{tr}$ of task $\mathcal{T}_{t}$, CM selects a subset by maximizing the coverage of the attribute/embedding space:
(2) $\mathcal{N}(v_{i})=\{v_{j}\mid\text{dist}(v_{i},v_{j})<d,\,\mathcal{Y}(v_{i})\neq\mathcal{Y}(v_{j})\},$
@@ 4.2. Diversity Sampling with semantic Consistency
(3) $\mathcal{C}_{r}=\bigcup_{v\in V_{\tau_{t}}}\left\{u\,\middle|\,(v,u)\in E_{r}\lor(u,v)\in E_{r}\right\},$
@@
(5) $V^{*}_{\tau}=\text{topk}_{v\in\mathcal{C}_{\tau}}\left(\pi(v),|\mathcal{B}_{\tau}|\right),$
@@
(6) $V^{\prime}=V_{\tau_{t}}\cup\bigcup_{\tau\neq\tau_{t}}V_{\tau},\quad E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\},$
@@
(7) $\mathcal{L}$ $=\mathcal{L}_{\mathcal{T}_{t}}(\mathcal{G}_{t},\theta)+\lambda_{er}\mathcal{L}_{er}(\mathcal{G}_{sub},\theta)$
@@
where $\hat{\mathcal{Y}}(v_{i})$ is the predicted label of node $v_{i}$, $\lambda_{er}$ modulates the contribution of buffered data in the overall loss.
4.3. Heterogeneity-aware Knowledge Distillation for Task Alignment
@@
Logit-level Distillation Inspired by previous distillation-based approaches (Tian et al., 2023), we transfer the soft knowledge from the teacher model (previous task classifier) to the student model (current task classifier), enabling the student model to learn both new knowledge from the current task and retained knowledge from past tasks. Mimicking teacher’s prediction results enables the student model to learn the secondary information from previous tasks that cannot be expressed by the experience replay data alone. Soft knowledge from teacher model is formulated as the predicted probability of the labels in the current task data:
(8) $P^{T}(z_{i},t)=\text{Softmax}\big(f_{T}(z_{i}),t\big)=\frac{\exp\left[f_{T}(z_{i})/t\right]}{\sum_{j}\exp\left[f_{T}(z_{j})/t\right]},$
where $z_{i}$ is the embedding of node $v_{i}$ in $\mathcal{V}_{i}^{tr}$, $f_{T}(z_{i})$ is the score logit of $z_{i}$ obtained from teacher model, and $t$ is the temperature index to soften the peaky softmax distribution (Hinton, 2015). Thus, the knowledge distillation loss of the teacher model and the student model is defined as follows:
(9) $\mathcal{L}_{\text{logit}}=\text{Mean}\big(t^{2}\sum_{l}^{N}\sum_{v_{i}\in\mathcal{V}_{i}^{tr}}P_{l}^{T}(z_{i},t)\log\frac{P_{l}^{T}(z_{i},t)}{P^{S}(z_{i},t)}\big),$
@@
(10) $\mathcal{L}_{\text{sem}}=\sum_{m=1}^{M}\|\boldsymbol{\alpha}_{P_{m}}^{(T)}-\boldsymbol{\alpha}_{P_{m}}^{(S)}\|_{2}$
where $\boldsymbol{\alpha}_{P_{m}}\in\mathbb{R}^{|V^{tr}|}$ is the normalized attention vector over all nodes for metapath $P_{m}$}. We provide a detailed discussion in Appendix A on how this module can be extended to arbitrary HGNNs beyond metapath-based architectures.
Algorithm 1 Details of the HERO Framework.
1:Input: Continual tasks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$; Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$; Learning rate $\alpha$ for the inner update; Learning rate $\beta$; Number of target type nodes in each class added to $\mathcal{B}$: e.
2:Output: Model $f_{\theta}$ which can mitigate catastrophic forgetting of preceding tasks.
3:Initialize model parameters $\theta$
4:while Continual Tasks $\mathcal{T}$ remains do
5:  Obtain current training set $V^{tr}_{t}$
6:  Obtain teacher models $\{f_{T}\}$ and Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$
7:  Select meta-learning examples $\mathcal{E}=Select(\mathcal{V}_{i}^{tr},e)$
8:  for $epoch=1$ to $E$ do
9:   Update the parameters $\theta$ using $\mathcal{E}$ via Equation 1
10:   Replay the experience via Equation 7
11:   Distill the soft knowledge and semantic information from teacher models via Equation 9 and 10
12:   Update the parameters of student model by optimizing Equation 12
13:  end for
14:  Select target type nodes via 2 and other types nodes via 5
15:  Add all sampled nodes to Buffer
16:  Add current task model to teacher models $\{f_{T}\}$
17:  $\mathcal{T}=\mathcal{T}\setminus\{\mathcal{T}_{i}\}$
18:end while
19:Return Model $f_{\theta}$
@@
(11) $\mathcal{L}_{kd}=\lambda_{\text{logit}}\mathcal{L}_{\text{logit}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}},$
@@
(12) $\mathcal{L}_{joint}=\mathcal{L}_{\mathcal{T}_{t}}+\lambda_{er}\mathcal{L}_{er}+\lambda_{kd}\mathcal{L}_{kd},$
@@
(13) $\theta^{*}=\theta-\beta\nabla_{\theta}\mathcal{L}_{joint}(\theta),$
where $\beta$ is learning rate. By minimizing $\mathcal{L}_{joint}$, parameters $\theta$ of HGNN model are optimized for downstream node classification task. As the number of tasks grows, the teacher models would increase linearly. For scalability, we adapt a sliding window strategy that only keeps the latest three teacher models for knowledge distillation. The details of the HERO framework is provided in 1.
Table 1. Performance comparison with different HGNN backbones on three benchmark datasets. The symbol $\uparrow$ ($\downarrow$) means higher (lower) is better. The best results are highlighted in bold, while the second best results are underlined. ”OOM” means that the model runs out of memory on large graphs.
@@
EWC (Kirkpatrick et al., 2017) 85.4 $\pm$ 5.5 21.5 $\pm$ 10.6 69.6 $\pm$ 1.2 25.8 $\pm$ 2.4 58.3 $\pm$ 2.4 18.4 $\pm$ 4.2
MAS (Aljundi et al., 2018) 91.1 $\pm$ 0.8 8.3 $\pm$ 1.5 72.0 $\pm$ 2.3 20.5 $\pm$ 3.9 59.0 $\pm$ 1.0 14.8 $\pm$ 0.9
TWP (Liu et al., 2021) 90.5 $\pm$ 1.3 10.1 $\pm$ 2.6 74.8 $\pm$ 2.7 14.5 $\pm$ 4.4 60.9 $\pm$ 4.3 10.8 $\pm$ 6.9
ER-GNN (Zhou and Cao, 2021) 89.7 $\pm$ 2.9 13.1 $\pm$ 6.1 74.9 $\pm$ 2.3 12.7 $\pm$ 4.2 56.5 $\pm$ 1.3 19.6 $\pm$ 0.9
MetaCLGraph (Unal et al., 2023) 89.9 $\pm$ 0.9 11.4 $\pm$ 2.3 74.8 $\pm$ 3.2 14.8 $\pm$ 5.7 59.4 $\pm$ 2.5 13.0 $\pm$ 4.1
FTF-ER (Pang et al., 2024) 90.8 $\pm$ 1.4 9.8 $\pm$ 2.4 72.0 $\pm$ 4.2 19.7 $\pm$ 6.4 58.1 $\pm$ 0.7 14.6 $\pm$ 0.6
@@
EWC (Kirkpatrick et al., 2017) 89.1 $\pm$ 1.2 10.7 $\pm$ 1.8 77.3 $\pm$ 0.4 12.5 $\pm$ 0.2 69.6 $\pm$ 3.0 14.7 $\pm$ 2.9
MAS (Aljundi et al., 2018) 89.8 $\pm$ 2.5 9.0 $\pm$ 3.0 77.5 $\pm$ 1.2 11.8 $\pm$ 2.6 70.4 $\pm$ 3.1 14.8 $\pm$ 2.0
TWP (Liu et al., 2021) 89.1 $\pm$ 2.1 10.3 $\pm$ 1.1 77.7 $\pm$ 0.6 11.1 $\pm$ 1.6 69.6 $\pm$ 2.0 15.1 $\pm$ 0.9
ER-GNN (Zhou and Cao, 2021) 90.7 $\pm$ 1.5 3.5 $\pm$ 0.4 77.3 $\pm$ 2.2 9.0 $\pm$ 3.6 69.5 $\pm$ 1.4 13.6 $\pm$ 1.7
MetaCLGraph (Unal et al., 2023) 90.0 $\pm$ 0.7 10.9 $\pm$ 1.4 76.1 $\pm$ 0.3 12.2 $\pm$ 1.8 69.1 $\pm$ 2.2 16.3 $\pm$ 2.5
FTF-ER (Pang et al., 2024) 89.4 $\pm$ 2.2 5.5 $\pm$ 1.9 76.1 $\pm$ 1.4 8.9 $\pm$ 3.5 69.7 $\pm$ 2.8 13.0 $\pm$ 2.5
@@
EWC (Kirkpatrick et al., 2017) 91.0 $\pm$ 2.3 10.9 $\pm$ 4.5 73.4 $\pm$ 1.6 16.7 $\pm$ 2.5 OOM OOM
MAS (Aljundi et al., 2018) 92.1 $\pm$ 2.1 7.8 $\pm$ 5.3 74.3 $\pm$ 2.0 14.2 $\pm$ 3.7 OOM OOM
TWP (Liu et al., 2021) 91.6 $\pm$ 2.1 7.4 $\pm$ 5.0 74.0 $\pm$ 0.8 13.7 $\pm$ 3.3 OOM OOM
ER-GNN (Zhou and Cao, 2021) 90.0 $\pm$ 2.1 12.4 $\pm$ 4.3 75.3 $\pm$ 0.9 10.3 $\pm$ 1.5 OOM OOM
MetaCLGraph (Unal et al., 2023) 92.3 $\pm$ 0.4 2.1 $\pm$ 0.7 75.6 $\pm$ 1.0 11.8 $\pm$ 1.3 OOM OOM
FTF-ER (Pang et al., 2024) 92.8 $\pm$ 0.8 1.9 $\pm$ 2.3 73.0 $\pm$ 0.6 15.8 $\pm$ 1.4 OOM OOM
@@
5. Experiments
We conducted experiments on four widely used web graph benchmarks: DBLP (Lv et al., 2021), IMDB (Lv et al., 2021), Freebase (Lv et al., 2021) and Yelp (Yang et al., 2020) to evaluate the performance of HERO. We adopt three widely used HGNN backbones in our experiments: HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), and HGT (Hu et al., 2020). See Appendix  B.1 for the details of the datasets and  B.4 for experiment setup.
Baselines
To evaluate the effectiveness of our proposed method, we select strong baselines from the Continual Graph Learning Benchmark (CGLB) (Zhang et al., 2022a), including Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), Memory Aware Synapses (MAS) (Aljundi et al., 2018), Experience Replay GNN (ER-GNN) (Zhou and Cao, 2021), and Topology-aware Weight Preserving (TWP) (Liu et al., 2021). EWC and MAS were previously not designed for graphs, so we train a HGNN on new tasks but ignore the graph structure when applying continual learning strategies. Additionally, we include two experience-replay-based baselines: MetaCLGraph (Unal et al., 2023) and FTF-ER (Pang et al., 2024). Furthermore, we use Finetune method (without any continual learning techniques) as a lower bound, and Joint Train (which allows access to previous data during training) as an approximate upper bound (Caruana, 1997).
Metrics
In our experiments, Average Performance (AP) and Average Forgetting (AF) (Lopez-Paz and Ranzato, 2017), are used to measure the performance on test sets. AP and AF are defined as $\text{AP}=\frac{1}{T}\sum_{t=1}^{T}a_{T,t},\,\,\,\,\text{AF}=\frac{1}{T-1}\sum_{t=1}^{T-1}\left(a_{T,t}-a_{t,t}\right),$ where $T$ is the total number of tasks and $a_{i,j}$} is the accuracy of the model on the test set of task $j$ after it is trained on task $i$.